A horizontal force of 21 N is applied to a 20 kg block lying on a horizontal surface. Find the acceleration of the block, if the coefficient or kinetic friction is 0.10.


1
Expert's answer
2021-07-07T08:38:55-0400

Let's apply the Newton's Second Law of Motion:


"F_{h}-F_{fr}=ma,""F_{h}-\mu_kmg=ma,""a=\dfrac{F_{h}-\mu_kmg}{m},""a=\dfrac{21\ N-0.1\cdot20\ kg\cdot9.8\ \dfrac{m}{s^2}}{20\ kg}=0.07\ \dfrac{m}{s^2}."